OpenStudy (anonymous):

Solve the inequality. –6b > 42 or 4b > –4

OpenStudy (compassionate):

Treat it like an equation. -6b > 42 --- --- = (b < -7) That's your first one. -6 -6 Can you do the other one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

honestly

OpenStudy (compassionate):

Do you know how to solve equations?

OpenStudy (anonymous):

i have tried many times

OpenStudy (compassionate):

Would you like me too teach you?

OpenStudy (anonymous):

get lost with alphabets

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@Compassionate are you here

OpenStudy (compassionate):

Allow me to introduce myself, I am a Ph.D graduate and full time professor. Solving equations is really simply. All you have to do is remember a few things. Rule 1. Do the opposite when solving an equation. Rule 2. Solve for the variable last. (The variable is the letter.) So lets start... If I had something like: 2 + 2x = 10, can you find the variable? Yes! The variable is the, "x." because it is the letter in the equation. So now that we found our letter, what do you we now? Well, as the rules state we do the opposite. So we have the equation: 2 + (2x) = 10. The parenthesis say that we're going to do that part last, because you always solve for your variable last. What do we mean by doing the opposite? Well you see that, "2," in the equation? Well it's positive, right? It's positive two! So we are going to do the opposite and subtract 2. The equation now looks like this: -2 + 2 + 2x = 10 <--- This looks nasty. However, you see the we have negative two + positive two, which just equals 0. So we are now left with 0 + 2x = 10 And this is equal too: 2x = 10(This is because 0 has no value and is useless.) Now you HAVE to subtract two from the other side of the equation to maintain equality. You do this for all equations. Example: 2 + 1x = 3 (-2 )+ 2 + 1x = 3 (- 2) ---------------------- 1 + 1x = 2 (-1) + 1 + 2x = 2(-1) ---------------------- You can see that what we do too one side of the equation we must do to the other. It is how we maintain equality. Imagine a scale, and both sides are equal. If I take five pounds of the left side, than the scale is unequal. In order to maintain quality we have too take five pounds off the other side of the scale. That is why we call it an "equation" because both sides must be equal to each other. So after you do the opposite one one side, you must take that number and do the same thing to the other side. Now the only term remaining is our variable. You can see that our variable is right next to a number. When you have something like 4x, 2x, 1x, etc, etc, etc.. All that is saying is, "Four times x," "Two times x," "One times x," etc. So we have: 2x = 10 - 2 (Ten minus two is eight.) 2x = 8 Now that we know it is being multiplied, we have to do the opposite until only the variable is remaining. What is the opposite of multiplication? That's right! Division! So we divide 2 by itself and get 1 because 2 can go into itself once! 2 - = 1 2 Now we must also divide the other side of the equation by 2! 1x = 10 -- 2 10 can go into 2 how many times? That's right! five times! We now have: 1x = 5 <--- Correct! Here are some example problems and tips. Tips: The opposite of addition is subtraction The opposite of subtraction is addition The opposite of division is multiplication The opposite of multiplication is division. Example Problems: 2x + 5 = 27 implies: 2x + = 27 − 5 = 22 (We subtracted five from itself and got 0, then we did the same thing to the other side of t equation and got 22 because 27 - 5 = 22) I mplies: 2x = 22 (Two can go into itself once, and into 22 11 times.) -- -- 2 2 implies: 1x = 11. Example: Solve for x : 5 − 2x = −11 −2x = −11 − 5 (A minus, minus a minus = a positive.) {That is -11 -5} 2x = 16 x = 8 Tutorial requested by Daboss33